OpenStudy (anonymous):

If log(x + 3) – log(x – 5) = log 2, what is the value of x? A-x = 2 B-x = 13 C-x = 5 D-x = -30

rishavraj (rishavraj):

\[\log a - \log b = \log \frac{ a }{ b } \]

OpenStudy (anonymous):

so the answer is ?

OpenStudy (anonymous):

@rishavraj

rishavraj (rishavraj):

x = 13

rishavraj (rishavraj):

@bettylee